The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Standard Hours Standard Rate Standard Cost Motor tune-up 2.5 $25.00 $62.50 The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 50 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups: Labor rate variance $ 150 F Labor spending variance $ 200 U Required: 1. Determine the number of actual labor-hours spent on tune-ups during the week. 2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labour efficiency variance = 200+150 = 350 U

a) Labour efficiency variance = (Standard hour-actual hour)Standard rate

-350 = (50*2.5*25-25X)

-350 = 3125-25X

25X = 3475

X(actual hours) = 3475/25 = 139 Hours

Labour rate variance = (Standard rate-actual rate)actual hours

150 = (139*25-139X)

139X = 3325

X(actual rate per hour) = 3325/139 = 23.92 per hour